In the synthesis of ammonia, what is the percentage yield of ammonia when 100.kg of hydrogen reacts with 800.0kg of nitrogen to produce 400.0kg of ammonia?
Answer #1

6kg of hydrogen requires 28 kg of nitrogen to produce 34 kg of ammonia.

thus 100 kg of hydrogen requires (100/6)*28 kg =466.67 kg of nitrogen to produce (100/6)*34 kg=566.67 kg of ammonia

But 400 kg of ammonia was obtained.Thus percentage yield is (400/566.67)*100 %=82.35 %
Answer #2

Just put values in this example

Calculating Percentage Yield

112g of nitrogen gas reacts with hydrogen gas to produce 40.8g of ammonia gas according to the equation given below:

N2(g) + 3H2(g)eqlarrow.gif 2NH3(g)

Calculate the percentage yield of ammonia.

  1. Actual yield of ammonia (NH3) = 40.8g
  2. Theoretical yield of ammonia (NH3) is calculated using the equation:

    moles = mass